Recent Log Entries

COMMENT ON Test Dominion Automatic Voltage Regulators

Log Number: 3876803Logbook(s): Body: 

Paul Powers, Todd Kujawa, and the local Dominion VA Power technical staff have been on site and waiting for the Richmond Operations Center to orchestrate today's tests.

The plan remains to conduct the test today, but it is clearly going to be later than the planned 9am to 11am.

CREW CHIEF SIGN IN

Log Number: 3876800Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 05-12-21
Time : 10:18:28
Day : Wednesday
Shift : Day

Crew Chief : Merz
Program Deputy: Aiken
Team : D
Arm : RADCON
Designated SSO: see Schedule
Operators : see Schedule

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

iocinjharp2 restored

Log Number: 3876799Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2021-05-12 10:12:53 The output from the myRestore execution was as follows: iocinjharp2 pcharp@2021-05-12 10:12 6(6) restored systemEnum@2021-05-12 10:12 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocinjharp2 (success) Procedure return status: Success

iocse4 restored

Log Number: 3876794Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2021-05-12 09:56:04 The output from the myRestore execution was as follows: iocse4 system@2021-05-12 09:56 2(2) restored seebpm@2021-05-12 09:56 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse4 (success) Procedure return status: Success

IPMAL01 has been enabled

Log Number: 3876793Logbook(s): Body: RF module was replaced for this bpm and passed calibration. Looks like ioc was restored from time bpm was disabled. Re-enabling bpm.

IPM8L01 has been enabled

Log Number: 3876792Logbook(s): Body: RF module was replaced and bpm had passed calibration. Looks like IOC was restored from time when bpm was disabled. Re-enabling bpm.

IPM6L01 has been enabled

Log Number: 3876791Logbook(s): Body: RF module had been replaced and bpm passed calibration. Looks like ioc was restored when bpm was disabled. Re-enabling bpm.

IPM4L01 has been enabled

Log Number: 3876790Logbook(s): Body: RF module was replaced for this bpm and passed calibration. Looks like the ioc was restored from a time when this bpm was disabled. Re-enbaling the bpm.

IPM2L01 has been enabled

Log Number: 3876789Logbook(s): Body: The RF module was replaced and had been passing calibration. Looks like the ioc was restored from a period that had bpm disabled. Re-enabling bpm.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3876788Logbook(s): Body: 

Day #158, Week 32 Day 3, Plans:

Tunnel --
* 830-9a: Cryo new employee tour (Wilson, +2)
* 0I02 harp controls checkout, followed by 0I02 anti-collision circuit (harp-YAG) testing (Dela Cruz)
* ItoV and picoammeter system checks (Francis)
* Connecting Wien HV cables to electrodes, verifying polarity interlocks, and beginning to apply to HV while monitoring vacuum (Dela Cruz, Hernandez-Garcia)
* Ion pump noise investigation (Stutzman)
* Valve airline upgrades from poly to stainless steel (Adderley)
* Laser hut work (Zhang)
* QE scan of freshly activated cathode (Zhang, Grames)
* Afternoon, conti

iocinjharp2 restored

Log Number: 3876786Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2021-05-12 09:11:56 The output from the myRestore execution was as follows: iocinjharp2 pcharp@2021-05-12 09:11 6(6) restored systemEnum@2021-05-12 09:11 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocinjharp2 (success) Procedure return status: Success

COMMENT ON Install and Commission Wien filter Power Supplies

Log Number: 3876782Logbook(s): Body: 

Late Entry: Troubleshooting of the Injector Wein filters revealed that the low resistance and inductive characteristics of the magnets are contributing to the sagging of the input AC voltage. Additional loading will need to be added.

APPROVED Breaker exercise buildings 04, 21, 07, 37

Log Number: 3876774Logbook(s): Body: 

The breaker exercise will occur in Bldg. 21 (E1 srvc), 7 (Ex. Stair 1), 37 (Ex. Stair 2), and 4 (Ex. Stair 4) this coming Friday (May 14) at 0830.

This task will consist of many brief power outages occurring intermittently to area systems and devices.

The following systems and devices will be affected:

VESDA
Fire Alarms
Rad monitors
Dehumidifiers (Tunnel/Outside)
Trim racks
PSS racks
iocea1/ea1mag

Please consider powering down your system if this exercise may present an issue. If there are any questions, please contact me (x7296)

Approved

IOC: mc1 Crash Recovery with settings from hour normal.2020-03-24.07.09.29

Log Number: 3876772Logbook(s): Tags: Body: 

keyword=IOC: mc1 Crash Recovery with settings from hour normal.2020-03-24.07.09.29
Any error messages from restore are below:

-> Preparing to restore MastOsc parameters -- press return to continue \n

-> Preparing to restore bpm parameters -- press return to continue \n
-> Preparing to restore Demand Power settings -- press return to continue \n

-> Preparing to restore OPSignals parameters -- press return to continue \n

-> Preparing to restore lock parameters -- press return to continue \n

-> Preparing to restore overview parameters -- press return to continue \n

-> Pr

iocse5 restored

Log Number: 3876771Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2021-05-09 06:29:25 The output from the myRestore execution was as follows: iocse5 system@2021-05-09 06:29 2(2) restored seebpm@2021-05-09 06:29 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse5 (success) Procedure return status: Success

iocse5 restored

Log Number: 3876770Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2021-05-12 06:22:32 The output from the myRestore execution was as follows: iocse5 system@2021-05-12 06:22 2(2) restored seebpm@2021-05-12 06:22 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse5 (success) Procedure return status: Success

COMMENT ON Hall A Corrector Magnet Re-Work

Log Number: 3876768Logbook(s): Body: 

Mike Beck
Tue 5/11/2021 3:20 PM
All,

We have completed all QA on the Hall-A blue dipole. To include
leak check, flow testing, hipot testing and core field magnet
measurements @ 100/200/300A. All pass acceptance testing
requirements.

We transferred the dipole to the Install shop this afternoon for
fab/brazing work. When completed, Install will bring the magnet
back to the Magnet Shop for temp switch install and electrical
testing of klixon safety system.

Mike

Pages